OpenStudy (anonymous):

8x9y=24.5 7y-2x=-8.5 solve by the elimination method

OpenStudy (anonymous):

I assume the first equation is missing an addition or subtraction symbol.

OpenStudy (anonymous):

ya its suppose to be 8x-9y=24 sorry

OpenStudy (anonymous):

8x-9y=24.5 -8x+28y=-34 multiply each term by 4 and add 19x=-9.5 x=-1/2 is it -24 or -24.5?

OpenStudy (anonymous):

y = 10/37 x = 3.3

OpenStudy (anonymous):

wait ..this is assuming 24

OpenStudy (anonymous):

just 24.5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it says that it is 8x-9y=24.5 7y-2x=-8.5

OpenStudy (anonymous):

and it says use elimination method

OpenStudy (anonymous):

y = -9.5/37 x = 3.35

OpenStudy (anonymous):

how did you get that answer i am trying to learn thats why I am asking

OpenStudy (anonymous):

multiply the 2nd eq(eq2) by 4 you get a -8x term in it on adding eq1 and eq2 that term gets cancelled leaving only 37y = 24.5 -4*8.5...this gives you y.Then take the value of y and put it in any eq this will give an eq purely in x. p.s.Sorry for not posting an explanation with my earlier answer...hope this helped

OpenStudy (anonymous):

HOW WOULD I WRITE THIS AS A SOLUTION